OpenStudy (anonymous):

Question 1 The expression 11x2 - 12x - 10 has ___ terms and a constant of ___. Answer 12, 11 3, -10 3, 10 11, 12 *I will choose B

OpenStudy (anonymous):

@HourglassMage

OpenStudy (anonymous):

B, yes

OpenStudy (anonymous):

Question 2 Combine like terms in the following expression: 9x + x. Answer 9x2 x + 7 x 10x 8x *C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Question 3 Use the distributive property to write the expression –6(–3x – 5) without parentheses. Answer 18x + 30 -18x – 30 18x – 5 -18x - 5 *A

OpenStudy (anonymous):

Erm...difficult to read the expression..can you type it?

OpenStudy (anonymous):

Okay, yes, it appeared

OpenStudy (anonymous):

You are correct :)

OpenStudy (anonymous):

Question 4 Evaluate for a = 5 and b = 2. Answer -25 85 33 18 *C

OpenStudy (anonymous):

can you type the equation please?

OpenStudy (anonymous):

\[4 + a ^{2} -4 \times (b-3)\]

OpenStudy (anonymous):

Question 5 Translate the following statement into an appropriate algebraic equation: The number of dogs (d) in the animal shelter is twice the sum of 4 and the number of cats (c). Answer d=2c+4 d= 2+4+c d= 2c - 4 d= 2(4+c)

OpenStudy (anonymous):

You were correct with Question 4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

For question 5 I chose D

OpenStudy (anonymous):

Question 6 Which algebraic expression represents the statement: the product of a number and twelve. Answer n + 12 n – 12 12n *c

OpenStudy (anonymous):

Question 7 Which equation best expresses the revenue for selling n books at $12.50 each? Answer r+n=12.5 r=n+12.5 r= 12.5 / n r= 12.5n *D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Six would be n + 12

OpenStudy (anonymous):

words that signal addition are 'and' words that signal multiplication are 'by'

OpenStudy (anonymous):

there are others, they're not coming to mind right now

OpenStudy (anonymous):

But question 6 ask us about Which algebraic expression represents the statement: the product of a number and twelve.===> so I think talking about product will be multiplication right? That means 12n

OpenStudy (anonymous):

oh...brain fart yes you're right missed that entirely

OpenStudy (anonymous):

You're right. :) Sorry, my mistake.

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

Question 9 What is the value of x in the following equation: 5=25 + x ? Answer 30 20 -20 -30

OpenStudy (anonymous):

Yes, you're right :)

OpenStudy (anonymous):

May I ask is number 7 correct?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

11 and 12 are right

OpenStudy (anonymous):

13 and 14 are right

OpenStudy (anonymous):

yes...thank :)

OpenStudy (anonymous):

15 - 18 are correct

OpenStudy (anonymous):

How about number 10, 7,

OpenStudy (anonymous):

Yes, both are right

OpenStudy (anonymous):

actually, hold on, you didn't post Question 10

OpenStudy (anonymous):

Seven is correct, but I'll need to see Question 10 to check it

OpenStudy (anonymous):

23 - 25 are correct

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Yeah Question10 was super easy like 14-10 =? i can answer that right?

OpenStudy (anonymous):

100% sis......*speechless*

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Woohoo!! :D That was all you though, no changes needed :D